Five Measures of Solvency or Profitability

The balance sheet for Garcon Inc. at the end of the current fiscal year indicated the following:

Bonds payable, 8% $900,000
Preferred $5 stock, $100 par $228,000
Common stock, $8 par $319,200.00

Income before income tax was $201,600, and income taxes were $30,600 for the current year. Cash dividends paid on common stock during the current year totaled $33,516. The common stock was selling for $28 per share at the end of the year.

Determine each of the following. Round answers to one decimal place, except for dollar amounts which should be rounded to the nearest whole cent. Use the rounded answers for subsequent requirements, if required.

a. Times interest earned ratio times
b. Earnings per share on common stock $
c. Price-earnings ratio
d. Dividends per share of common stock $
e. Dividend yield
Answer #1

a. Times interest earned ratio => EBIT/Interest

ie., = (201600+72000)/72000 = 3.8 Times

b. EPS on common stock = (EBT - Tax - Pref. dividend)/No of common stock

ie., (201600 - 30600 - (NA))/(319200/8) => 4.29 $ per share.

c. PE Ratio => Market value per share/EPS ie., 28/4.29 = 6.5 Times

d. Dividend per share => $ 33516/39900 = $ 0.84/ share

e. Dividend yield = Dividend per share/Current market price*100 => 0.84/28*100 = 3%